Find the exact value of sin θ/2 if cos θ = 2/3 and 270º< θ<360º.
Thank You

Answer by fractalier(2101)

Let q be theta...
sin q/2 = ± sqrt((1 - cos q) / 2)
sin q/2 = ± sqrt((1 - 2/3) / 2)
sin q/2 = ± sqrt(1/6) = ± sqrt(6) / 6
In the 4th quadrant, sin is negative, but in the 2nd quadrant, where q/2 is, the sin is positve, thus
sin q/2 = sqrt(6) / 6